### v1

@@ 5.1 Case (A)
It is consistent that ${\mathbb{P}}(\omega^{\kappa+1})$ collapses $2^{\kappa}$ to $\omega_{1}$, although $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})<(2^{\kappa})^{+}$ and ${\mathbb{P}}(\omega^{\kappa})$ does not collapse $2^{\kappa}$ to $\omega$. First, concerning claim (c) of Theorem 5.2(c) we note that the equality ${\mathfrak{h}}=\omega_{1}$ holds in Cohen extensions of models of GCH; then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. If $V^{\prime}$ is a Cohen model for ${\mathfrak{c}}=\omega_{5}$ obtained from a model $V$ of GCH, then in $V^{\prime}$ we have ${\mathfrak{h}}=\omega_{1}$ and, by (9), $2^{\omega_{1}}=(\max(\omega_{5},2)^{\omega_{1}})^{V}=(\omega_{5}^{\omega_{1}})^{V}=\omega_{5}$. By Theorem 4.10(A) and Fact 2.8(e) we have $1_{{\mathbb{P}}(\omega^{\omega_{1}+1})}\Vdash|2^{\omega_{1}}|=\omega_{1}$ although (see [5]) $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}<\omega_{6}=(2^{\omega_{1}})^{+}$ and forcing by ${\mathbb{P}}(\omega^{\omega_{1}})$ does not collapse $2^{\omega_{1}}$ to $\omega$.
Countable cofinality
Here we consider the poset ${\mathbb{P}}(\omega^{\delta})$ in the particular case, when $\delta=\mu$ is a cardinal and $\mu>\mathop{\rm cf}\nolimits(\mu)=\omega$. As in the previous subsection we are interested when $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$ (and, hence, the collapse by ${\mathbb{P}}(\omega^{\delta})$ is maximal) and we analyze the relationship between the forcing ${\mathbb{P}}(\omega^{\delta})$ and its first iteration, equivalent to $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. By Fact 2.12 we have $\omega^{\delta}=\omega^{\mu}=\mu$; thus ${\mathbb{P}}(\omega^{\delta})={\mathbb{P}}(\mu)=[\mu]^{\mu}$ and by Fact 2.2(f) we have
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))=\mathop{\rm ro}\nolimits({\mathcal{P}}_{\mu}).$ (45)
Regarding the assumptions of Fact 2.6, the poset ${\mathcal{P}}_{\mu}$ is separative, atomless, $\omega_{1}$-closed and $|{\mathcal{P}}_{\mu}|=2^{\mu}$ (Fact 2.7(a) and Theorem 4.10) we have $1_{{\mathcal{P}}_{\mu}}\Vdash|\mu^{\omega}|=\omega_{1}$ (Fact 2.8(f)), and $(2^{\mu})^{<\omega_{1}}=2^{\mu\omega}=2^{\mu}$. Thus, since in this context forcing equivalence of posets implies isomorphism of their Boolean completions, by (45) we have
$\mu^{\omega}=2^{\mu}\Rightarrow\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu}).$ (46)
We note that, by König’s theorem,
$\mu^{+}\leq\mu^{\omega}=\mu^{\mathop{\rm cf}\nolimits(\mu)}\leq 2^{\mu},$ (47)
so, ${\mathbb{P}}(\omega^{\delta})$ collapses $\mu^{+}$ to $\omega_{1}$. Since $\omega_{1}\leq{\mathfrak{h}}\leq{\mathfrak{c}}=2^{\omega}\leq\mu^{\omega}$ and $1_{{\mathcal{P}}_{\mu}}\Vdash|\mu^{\omega}|=\omega_{1}$, ${\mathbb{P}}(\omega^{\delta})$ collapses ${\mathfrak{c}}$ (and, hence, ${\mathfrak{h}}$) to $\omega_{1}$ as well.
Theorem 5.5
If $\delta=\mu$ is a cardinal and $\mu>\mathop{\rm cf}\nolimits(\mu)=\omega$, then each of the following conditions implies that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$
1. $2^{\mu}=\mu^{+}$
2. $2^{<\mu}=\mu$
3. $\mathrm{MA}+\mu<{\mathfrak{c}}$
4. $V$ is a Cohen model for ${\mathfrak{c}}=\kappa$ obtained from a model of GCH in which $\kappa$ is regular and $\mu<\kappa$.
In case 4 we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ while this is false in cases 1 if e.g., ${\mathfrak{c}}=\omega_{1}$, and in 3.
Proof. 1. If $2^{\mu}=\mu^{+}$, then by (47) we have $\mu^{\omega}=2^{\mu}$ and if, for example, ${\mathfrak{c}}=\omega_{1}$, then by (46), $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})\neq\mathop{\rm Col}\nolimits(\omega_{1},\omega_{1})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$.
2. Since for a singular cardinal $\mu$ we have $2^{\mu}\in\{2^{<\mu}\mu^{\mathop{\rm cf}\nolimits(\mu)},(2^{<\mu})^{\mathop{\rm cf}\nolimits(2^{<\mu})}\}$, if $\mu$ is a strong limit cardinal, that is, $2^{<\mu}=\mu$, then $\mu^{\omega}=2^{\mu}$ and by (46) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$.
3. If $V\models\mathrm{MA}+{\mathfrak{c}}>\mu$, then, since $\mu<{\mathfrak{c}}$ by (47) we have ${\mathfrak{c}}=2^{\omega}\leq\mu^{\omega}\leq 2^{\mu}={\mathfrak{c}}$; so, $\mu^{\omega}=2^{\mu}$ and by (46) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$. Here the first iteration $\langle[\omega]^{\omega},\subset^{*}\rangle$ does not collapse cardinals (because ${\mathfrak{h}}={\mathfrak{c}}$) and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})\neq\mathop{\rm Col}\nolimits({\mathfrak{c}},{\mathfrak{c}})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$.
4. Let $V$ be a model of GCH and, in $V$, let $\kappa>\mu$ be a regular cardinal. If $V^{\prime}=V[G]$ is a Cohen model for ${\mathfrak{c}}=\kappa$, obtained by ${\mathbb{P}}=\mathop{\mathrm{Fn}}\nolimits(\kappa,2)$, then by (9) in $V^{\prime}$ we have $2^{\mu}=(\max(\kappa,2)^{\mu})^{V}=(\kappa^{\mu})^{V}=\kappa$ and $\mu^{\omega}=(\max(\kappa,\mu)^{\omega})^{V}=(\kappa^{\omega})^{V}=\kappa$. So, by (46) in $V^{\prime}$ we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$. Since in $V^{\prime}$ we have ${\mathfrak{h}}=\omega_{1}$ (see [11], p. 256) and, by Fact 2.8(e), forcing by $\langle[\omega]^{\omega},\subset^{*}\rangle$ collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$, and ${\mathfrak{c}}^{<\omega_{1}}={\mathfrak{c}}^{\omega}={\mathfrak{c}}$, by Fact 2.6 we have $\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)=\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits\langle[\omega]^{\omega},\subset^{*}\rangle)\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$. So, the two-step iteration is forcing equivalent to $P(\omega)/\mathop{\rm Fin}\nolimits$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})=\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$. $\Box$
Remark 5.6
By Fact 2.8(f) we have $\mathop{\rm Col}\nolimits(\omega_{1},\mu^{\omega})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\mu})\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))$; but the equality $\mathop{\rm Col}\nolimits(\omega_{1},\mu^{\omega})\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))$ is not a theorem of ZFC. Baumgartner [5] constructed a model of ZFC in which ${\mathcal{P}}_{\mu}$ contains an adf of size $2^{\mu}>\mu^{\omega}$ and, hence $\pi(\mathop{\rm ro}\nolimits({\mathcal{P}}_{\mu}))=2^{\mu}>\mu^{\omega}=\pi(\mathop{\rm Col}\nolimits(\omega_{1},\mu^{\omega}))$.
Note that in 1 and 2 of Theorem 5.5 we have ${\mathfrak{c}}<\mu<\mu^{\omega}=2^{\mu}$ and that in 3 and 4 we have $\mu<{\mathfrak{c}}=\mu^{\omega}=2^{\mu}$. Generally, if $\mu<{\mathfrak{c}}$, then $\mu^{\omega}\leq{\mathfrak{c}}^{\omega}=2^{\omega}\leq\mu^{\omega}$ and, hence, $\mu^{\omega}={\mathfrak{c}}$. But $\mu<{\mathfrak{c}}=\mu^{\omega}<2^{\mu}$ is consistent too. Let, in $V$, GCH hold let $S:=\{\aleph_{n}:n<\omega\}$, let $E:S\rightarrow\mathop{\mathrm{Card}}\nolimits$ be the Easton index function defined by $E(\aleph_{n})=\aleph_{\omega+n+2}$, and let ${\mathbb{P}}_{E}$ be the corresponding Easton forcing. Then (see [11], p. 265) in $V[G]$ we have $2^{\aleph_{n}}=\aleph_{\omega+n+2}$, for $n\in\omega$; and $2^{\aleph_{\omega}}=\aleph_{\omega+\omega+1}$. Thus taking $\mu=\aleph_{\omega}$ we have ${\mathfrak{c}}=\aleph_{\omega+2}=\mu^{\omega}<2^{\mu}=\aleph_{\omega+\omega+1}$. Thus $\mu^{+}<\mu^{\mathop{\rm cf}\nolimits(\mu)}<2^{\mu}$.
5.2 Cases (D) and (E)
Using Fact 2.5 and Theorem 4.10 it is easy to prove
Theorem 5.7
If (D) or (E) holds, $|\delta|=\mathop{\rm cf}\nolimits(\delta)=\kappa>\omega$ and $2^{\kappa}=\kappa^{+}$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$ and, in addition,
  • (a)
    If (D) holds, then $\delta$ satisfies one of the following conditions
@@ 5.2 Cases (D) and (E)
  • (a)
    If (E) holds, then $\delta$ satisfies one of the following conditions
@@
      where $\kappa=\mathop{\rm cf}\nolimits(\gamma)\leq\gamma<\kappa^{+}$ and ($\theta=0$ or $\kappa^{\gamma}\leq\theta<\kappa^{+}$);
If $\delta=\kappa>\omega$ is a regular cardinal we have case (D) and $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))={\mathcal{P}}_{\kappa}$. If $2^{\kappa}=\kappa^{+}$, then by Theorem 5.7 $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$.
@@
It is consistent that $\kappa\in\mathop{\mathrm{Reg}}\nolimits$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$. Baumgartner [5] constructed a model of ZFC in which, for $\kappa=\omega_{1}$, $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=\kappa^{++}<2^{\kappa}$. Thus ${\mathcal{P}}_{\kappa}$ collapses $\kappa^{+}=\omega_{2}$ to $\omega$ and the cardinals $\geq\kappa^{++}$ remain cardinals in extensions by ${\mathcal{P}}_{\kappa}$. So, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$.
In fact, if $\omega_{1}\leq\kappa\in\mathop{\mathrm{Reg}}\nolimits$, then by Fact 2.8(c),
$\forall\lambda\in[\omega,\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa}))\;\;1_{{\mathcal{P}}_{\kappa}}\Vdash|\check{\lambda}|=\check{\omega}\;\mbox{ and }\;|((\kappa^{+})^{V})\check{\;}|=\check{\omega}\;\mbox{ and }\;\omega_{1}=(\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})^{V})\check{\;}.$ (48)
By (48) Fact 2.7(a), and Fact 2.6 we have
$\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}\Leftrightarrow 1_{{\mathcal{P}}_{\kappa}}\Vdash|((2^{\kappa})^{V})\check{\;}|=\check{\omega}\Leftrightarrow\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa}).$ (49)
Singular cardinals
Let $\mathop{\mathrm{Card}}\nolimits\ni\delta>\kappa=\mathop{\rm cf}\nolimits(\delta)>\omega$ (the smallest such $\delta$ is $\aleph_{\omega_{1}}$). Since $\delta$ is right indecomposable and $\kappa\in\mathop{\mathrm{Reg}}\nolimits$, cases (B), (C) and (D) would give $\delta=\kappa$, which is false. So this is Case (E) and by Theorem 4.10 we have $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle[\delta]^{\delta},\subset^{*}\rangle$, which gives ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}{\mathcal{P}}_{\delta}\cong\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$. By Fact 2.7(b) ${\mathcal{P}}_{\delta}$ is $(\omega,\cdot,\delta^{+})$-nowhere distributive and by Fact 2.8(b)
$2^{\kappa}<\delta\Rightarrow\mathop{\rm Col}\nolimits(\omega,\delta^{+})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta}).$ (50)
If $2^{\delta}=\delta^{+}$ and $2^{\kappa}=\kappa^{+}$, then by Fact 2.9(a) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))=\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta})\cong\mathop{\rm Col}\nolimits(\omega,2^{\delta})$.
6 Ordinals of size $\omega_{1}$
@@
If $|\alpha|=\omega_{1}$, then $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed or $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$. More precisely, if $\alpha$ is presented in the Cantor normal form
$\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m,$ (51)
where $n,s_{1},\dots,s_{n}\in{\mathbb{N}}$, $1\leq\delta_{1}<\dots<\delta_{n}<\omega_{2}\,$ and $\,m\in\omega$, then
either $\forall i\leq n\;\Big{(}\mathop{\rm cf}\nolimits(\delta_{i})\leq\,\omega\,\;\lor\;\exists\theta\geq\omega_{1}\;(\mathop{\rm cf}\nolimits(\theta)=\omega\land\delta_{i}=\theta+\omega_{1})\Big{)},$ (52)
or $\exists i\leq n\;\Big{(}\mathop{\rm cf}\nolimits(\delta_{i})=\omega_{1}\;\land\;\Big{(}\exists\theta\geq\omega_{1}\;(\delta_{i}=\theta+\omega_{1})\Rightarrow\mathop{\rm cf}\nolimits(\theta)=\omega_{1}\Big{)}\Big{)}.$ (53)
(a) $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed and $((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})^{\sum_{i=1}^{n}s_{i}}\hookrightarrow\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ iff (52) holds;
(b) $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$ iff (53) holds;
@@
(c) If (53) holds and $\mathop{\rm cc}\nolimits(P(\omega_{1})/[\omega_{1}]^{<\omega_{1}})=(2^{\omega_{1}})^{+}$, then ${\mathbb{P}}(\alpha)\equiv\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$.
Proof. By Lemma 4.1, either (52) or (53) is true.
By Lemma 3.1, w.l.o.g. we suppose that $m=0$. Then defining $k:=\sum_{i\leq n}s_{i}$ we have $\alpha\cong\sum_{1\leq j\leq k}L_{j}:=L$, where $L_{j}\cong\omega^{\delta^{\prime}_{j}}$ and
$\langle\delta^{\prime}_{j}:1\leq j\leq k\rangle=\langle\delta_{n},\dots,\delta_{n},\delta_{n-1},\dots,\delta_{n-1},\dots,\delta_{1},\dots,\delta_{1}\rangle.$
($\delta_{i}$ appears $s_{i}$ many times). By Theorem 3.2 ${\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ and by Fact 2.2(d) and (e) $\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$.
(a) If (52) holds, then by Theorem 4.10 for each $j\in\{1,\dots,k\}$ the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ is $\sigma$-closed and $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\hookrightarrow\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$. By Facts 2.2(e) and 2.1, the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ is $\sigma$-closed.
(b) If (53) holds and $i_{0}<n$ witnesses that, then, by Theorem 4.10, $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta_{i_{0}}})\hookrightarrow_{c}\prod_{1\leq j\leq k}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$.
(c) By (b) we have $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$, which by Fact 2.3 implies that $\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)\equiv\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\ast\pi$. So, since $\mathop{\rm cc}\nolimits(P(\omega_{1})/[\omega_{1}]^{<\omega_{1}})=(2^{\omega_{1}})^{+}$, by Facts 2.2 and 2.8(a) and (c) the forcing $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ collapses $2^{\omega_{1}}$ to $\omega$ and, hence, $|\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)|=2^{\omega_{1}}$. Now, by Fact 2.6, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\equiv\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$ and by Fact 2.2(a) we have ${\mathbb{P}}(\alpha)\equiv\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$. $\Box$

### v2

@@ 5 The posets ${\mathbb{P}}(\omega^{\delta})$ and collapsing algebras
We regard Theorem 5.2 in the particular case when $|\delta|=\mathop{\rm cf}\nolimits(\delta)=:\kappa>\omega$ and $2^{\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)^{+}$. It is easy to check (using the Cantor normal form (4) for $\delta$) that if (D) holds, then $\delta$ satisfies one of the following conditions:
@@
On the other hand, if (E) holds, then $\delta$ satisfies one of the following conditions:
@@
    where $\kappa=\mathop{\rm cf}\nolimits(\gamma)\leq\gamma<\kappa^{+}$ and ($\theta=0$ or $\kappa^{\gamma}\leq\theta<\kappa^{+}$).
If $\delta=\kappa>\omega$ is a regular cardinal, then $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))={\mathcal{P}}_{\kappa}$. If $2^{<\kappa}=\kappa$ or $2^{\kappa}=\kappa^{+}$, then by Theorem 5.2 $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$.
@@
It is consistent that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$, for a regular cardinal $\kappa>\omega$. Baumgartner [2] constructed a model of ZFC in which $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}<2^{\omega_{1}}$. Thus ${\mathcal{P}}_{\omega_{1}}$ collapses $\omega_{2}$ to $\omega$ and the cardinals $>\omega_{2}$ are preserved in extensions by ${\mathcal{P}}_{\omega_{1}}$. So, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\omega_{1}})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$. In fact, if $\kappa>\omega$ is a regular cardinal, then by Fact 2.6(e), $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$ iff $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}$.
@@
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$, if $\kappa$ is a cardinal, $\kappa>2^{\mathop{\rm cf}\nolimits(\kappa)}>\mathop{\rm cf}\nolimits(\kappa)>\omega$ and $2^{\kappa}=\kappa^{+}$.
Proof. Since $\omega^{\kappa}=\kappa$ we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))={\mathcal{P}}_{\kappa}^{+}$ and by Fact 2.6(d) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))$ collapses $2^{\kappa}=\kappa^{+}$ to $\omega$. So, by Fact 5.1, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$. $\Box$
Collapse to $\omega_{1}$
Here we are looking for the ordinals $\delta$ such that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. First we show that if ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$ (examples of such $\delta$’s can be found in the previous paragraph), then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$, for all $n\in{\mathbb{N}}$, namely, the Boolean completion of ${\mathbb{P}}(\omega^{\theta})$ is a “constant function” on the interval $(\delta,\delta+\omega)$. We recall some facts from [10] and [13].
Fact 5.6
Let $\delta$ denote an ordinal, $\kappa$ a cardinal and ${\mathbb{B}}\,$ a Boolean algebra.
(a) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n}))\cong(\mathop{\rm rp}\nolimits^{n}(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}))^{+}$, for each $n\in{\mathbb{N}}$, whenever $\delta\geq 1$.
(b) $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega,\kappa)))\cong\mathop{\rm Col}\nolimits(\omega_{1},\kappa)$, for each $n\in{\mathbb{N}}$, whenever $\kappa^{\omega}=\kappa\geq{\mathfrak{c}}$.
(c) $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega_{1},\kappa)))\cong\mathop{\rm Col}\nolimits(\omega_{1},\kappa^{\omega})$, for each $n\in{\mathbb{N}}$, whenever $\kappa\geq 2$.
(d) $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm ro}\nolimits({\mathbb{B}})))\cong\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))$, for each $n\in{\mathbb{N}}$.
Proof. Statement (a) is a straightforward generalization of Lemma 3.2(d) from [10] (proved there for countable $\delta$). Statements (b), (c) and (d) follow from Theorems 6.9 and 5.4 and Fact 2.6 from [13]. $\Box$
Theorem 5.7
@@
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$, for all $n\in{\mathbb{N}}$, whenever $\delta\geq\omega_{1}$ and (a) or (b) holds, where
(a) ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$,
(b) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and collapses $2^{|\delta|}$ to $\omega_{1}$.
Proof. Defining $\lambda=\omega$, if (a) holds, and $\lambda=\omega_{1}$, if (b) holds, by Fact 5.1 we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$. Defining ${\mathbb{B}}:=P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$, by Fact 4.3 we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))={\mathbb{B}}^{+}$ and, hence, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))=\mathop{\rm ro}\nolimits({\mathbb{B}})$; thus, $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$. By Fact 5.6(a) we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta+n})\cong(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))^{+}$; so,
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta+n}))$$\cong$ $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))\quad\mbox{ (since $\mathop{\rm rp}\nolimits^{n}({\mathbb{B}})$ is a Boolean algebra)}$)
[…]
$\cong$ $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})))\quad\mbox{ (since $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$)}.$
Now, if (a) holds, then $\lambda=\omega$ and by Fact 5.6(b) we have $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. If (b) holds, then $\lambda=\omega_{1}$ and, by Fact 5.6(c), $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. $\Box$